### v1

# Tuning-Free Adaptive Style Incorporation for Structure-Consistent Text-Driven Style Transfer

Abstract
In this work, we target the task of text-driven style transfer in the context of text-to-image (T2I) diffusion models. The main challenge is consistent structure preservation while enabling effective style transfer effects. The past approaches in this field directly concatenate the content and style prompts for a prompt-level style injection, leading to unavoidable structure distortions. In this work, we propose a novel solution to the text-driven style transfer task, namely, Adaptive Style Incorporation (ASI), to achieve fine-grained feature-level style incorporation. It consists of the Siamese Cross-Attention (SiCA) to decouple the single-track cross-attention to a dual-track structure to obtain separate content and style features, and the Adaptive Content-Style Blending (AdaBlending) module to couple the content and style information from a structure-consistent manner. Experimentally, our method exhibits much better performance in both structure preservation and stylized effects.
Keywords:
Style transfer Diffusion model
00footnotetext: $*{}^{*}$qual contribution.
${}^{\dagger}$orrespoding author.
Figure 1: We propose Adaptive Style Incorporation (ASI), a tuning-free diffusion-based style transfer method that enables versatile text-guided stylization for the source image. Our stylized results exhibit high consistency to the structure and semantics of the source image, while significantly changing their image style following the style prompt.
1 Introduction
With the increasing prominence of text-to-image (T2I) [37, 35, 36] models featuring billion-parameter architectures trained on extensive text-image datasets, the prospect of translating our imaginative ideas into high-quality images through textual input becomes readily apparent. While these foundation models have ushered in a new era of creative expression, their powerful generation capability comes at the expense of limited controllability. Due to the text-image misalignment problem in the T2I models, even minor modifications to the textual prompt may result in unforeseen structure distortion in the output image. This poses a significant obstacle for artists and creators seeking precise artistic editing to their work.
In this study, we aim to investigate text-driven style transfer (see Fig. 1) within the context of T2I diffusion models. The major challenge in this field lies in preserving the original structure while producing effective stylized results. To resolve the aforementioned issue, numerous efforts have been made to incorporate additional conditional controls, such as structural guidance signals [49, 52] and editing masks [12, 2]. To be specific, BlendDiffusion [4, 2] and SpaText [3] train diffusion models that support a user-provided mask as input to achieve regional generation and editing. ControlNet [52] and Make-A-Scene [12] finetune diffusion models conditioned on the segmentation masks to ensure invariance of semantics in the editing images. However, these methods fall short of achieving fine-grained control over the structural layout and require extensive training on large-scale text-image pairs, thereby hindering intuitive and easily deployable text-driven style transfer.
Recently, several works [19, 43, 6, 1, 15] demonstrate that the attention features in diffusion models implicitly encode fine-grained structure information, and enables text-driven style transfer without the requirement for model tuning. This greatly reduces the cost of professional creation using the diffusion model. Among all these works, Prompt-to-Prompt (P2P) [19] and Plug-and-Play (PnP) [43] share the most similarity to ours methodologically. P2P proposes operating the cross-attention map from the source image diffusion process, allowing various editing capabilities while preserving the source image content. PnP proposes manipulating the self-attention map from the source image diffusion process, enabling better retention of fine-grained structural details.
However, the aforementioned approaches choose to directly concatenate the content prompt (representing the source image) and style prompt, in order to achieve prompt-level style injection. Nevertheless, due to the well-known text-image misalignment issue [23] in T2I models, the style prompt often contains non-style information, leading to unavoidable structure distortions and semantic drift in the stylized image for P2P and PnP.
Unlike previous studies that commonly inject styles on the coarse-grained prompt level, we propose Adaptive Style Incorporation (ASI) to incorporate the style information on the fine-grained feature level. In ASI, we first present the Siamese Cross-Attention (SiCA) that decouples the common single-track cross attention into a dual-track architecture, which utilizes the shared Query features to simultaneously query two sets of Key and Value features from the content prompt and style prompt individually, as depicted in Figure. 2 (b). Subsequently, we introduce a tuning-free Adaptive Content-Style Blending (AdaBlending) module to implement mask-guided style incorporation from a structure-consistent perspective. Thanks to the proposed ASI mechanism, our feature-level fine-grained style incorporation can maintain better structure consistency for the style transfer task.
Our contributions can be summarized as follows.
  • We propose a novel solution to the text-driven style transfer task, namely Adaptive Style Incorporation (ASI), to achieve feature-level fine-grained style incorporation. It preserves better image structure while enabling effective style transfer.
  • We propose the Siamese Cross-Attention (SiCA) and Adaptive Content-Style Blending (AdaBlending) modules to form the complete ASI mechanism. It requires no training/fine-tuning but rather leverages a pre-trained text-to-image diffusion model.
  • Experimental results have shown the effectiveness of our proposed ASI in both maintaining the structure-consistent layout and effectively embedding style attributes, whether applied to real images or generated images.
2 Related Work
2.1 Diffusion Models
The Diffusion Model [40] was initially employed to mitigate Gaussian noise applied continuously to training images and can be conceived as a series of denoising autoencoders. The advent of Denoising Diffusion Probabilistic Models (DDPM) [21] propelled the utilization of diffusion models for image generation into the mainstream. Subsequent advancements, exemplified by Denoising Diffusion Implicit Models (DDIM) [41], substantially accelerated sampling speeds while effecting directed generation from random noise to samples, facilitating subsequent research in image editing. Dhariwal and Nichol first introduced Classifier Guidance Diffusion [11], substantiating the superiority of diffusion models over GANs in image generation tasks. Classifier-Free Diffusion Guidance [22] then achieved a balance between sample quality and diversity during image generation. Building upon prior work, GLIDE [34] initiated using text as a conditional guide for image generation in a classifier-free manner. DALLE-2 [35], building on the foundation laid by GLIDE, incorporated CLIP as a guiding factor, bringing diffusion models into the public eye. Subsequently, the field of image generation based on diffusion models experienced a surge, with projects such as Imagen [37] and Stable Diffusion [36] emerging, prompting the gradual application of diffusion models in a broader range of tasks, including object detection [7], style transfer [46], image editing [29, 5, 8] and semantic segmentation [51]. With the release of ControlNet [52], the potential of diffusion models in various other domains is being further explored [10].
2.2 Style Transfer
Style transfer has garnered significant attention from researchers alongside the development of generative models. Early style transfer algorithms [13], such as AdaIN [24], primarily relied on optimization-based style transfer methods applied to the VGG architecture [39]. With the advent of Generative Adversarial Networks (GANs) [16], GANs have been employed in training style transfer models tailored to specific domains, such as ink paintings [17], landscape art [48], cartoons [45], and portraiture [50]. In addition, representative works such as CycleGAN [58], StarGAN [9], and StyleGAN [26] are devoted to the research of multi-domain style transfer. Currently, within the realm of style transfer research built upon the burgeoning diffusion model, there is a predominant focus on multi-domain style transfer studies. Research involving style transfer using diffusion models is akin to studies based on GANs, relying predominantly on loss functions for training. Representative works in this domain include StyleDiffusion [46], UCAST [55], CycleDiffusion [47], and InST [54], among others. A limited amount of research has focused on tuning-free style transfer. Some studies, such as StyleAligned [20], investigate style generation using images as style references. Plug-and-Play [43], on the other hand, explores style generation using text as a reference, and this study is the closest to our paper. However, these methods ignore injecting styles in a structurally consistent manner, thus disturbing the structural consistency between the stylized image and the source image. Compared to prompt-level coarse-grained one-shot style injection, we propose ASI which enables feature-level fine-grained style incorporation in a structure-consistent manner, achieving a good balance between stylization and structural preservation compared to other SOTA methods.
3 Preliminaries
The diffusion model [40] is a recently prominent probabilistic generative model, which accomplishes the image generation task by progressively predicting and removing Gaussian noise from a noisy image. The mathematical significance of the diffusion model is to simulate a distribution $p_{\theta}(x_{0})$ approximate to the data distribution $q(x_{0})$. Typically, diffusion models consist of two main processes: the forward process and the reverse denoising process. The forward process of noise injection is as follows,
[…]
$q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ (1)
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta_{t}}\epsilon,\epsilon\sim\mathcal{N}(0,\mathit{I}),$ (2)
[…]
where $t\in[0,T]$, $T$ is a preset maximum number of sampling steps, while the values of $\alpha_{t}$ are in a fixed variance schedule, $\alpha_{0}=1$ and $\lim_{t\to\infty}\alpha_{t}=0$. In practice, the forward noise injection process is a Markovian process, which means that given $x_{0}$, it is possible to obtain the noisy result for any number of steps directly,
[…]
$q(x_{t}|x_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0};(1-\bar{\alpha}_{t})\mathbf{I}),$ (3)
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\epsilon\sim\mathcal{N}(0,\mathit{I}),$ (4)
[…]
where $\alpha_{t}=1-\beta_{t}$ and $\hat{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
The reverse denoising process begins with $x_{T}$ as the starting point and involves sampling from the posteriors $q(x_{t-1}|x_{t})$ to perform continuous noise prediction. DDPM [21] sets the noise variance to a fixed value of $\sigma^{2}\mathbf{I}$, allowing the Gaussian transformation $p_{\theta}(x_{t-1}|x_{t})$ to be learned through a trainable $\mu_{\theta}(x_{t},t)$. Hence, the reverse denoising process can be represented as follows,
[…]
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t},t),\sigma^{2}_{t}\mathbf{I}),$ (5)
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\mu_{\theta}(x_{t},t))+\sigma_{t}\mathbf{z},$ (6)
[…]
where $z$ is Gaussian noise independent of $x_{t}$, and $\mu_{\theta}(x_{t},t)$ is a noise approximator $\epsilon_{\theta}(x_{t},t)$ optimized to mimic the forward noise $\epsilon$:
[…]
$f_{\theta}(x_{t},t)=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}},$ (8)
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}f_{\theta}(x_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}_{t}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\mathbf{z}}.$ (9)
[…]
One of the distinctive features of DDIM is that changing $\sigma_{t}$ in Eq. 9 can lead to different inverse processes. When $\sigma_{t}=0$, it essentially turns the mapping from latent to image into a fixed process, supporting perfect inversion of images. This is the basis for our acquisition of image latent.
Figure 2: Comparison between prompt-level coarse-grained style injection with our feature-level fine-grained style incorporation. (a) Due to the well-known issue of text-image misalignment in T2I models, directly concatenating the content prompt with the style prompt (i.e., prompt-level coarse-grained style injection) will introduce non-style information into the target image, resulting in unavoidable structural and semantic drift in the target image, such as the hair of the boy and the pattern on his clothes. (b) Compared to prompt-level coarse-grained style injection, we propose Adaptive Style Incorporation (ASI) to convert the existing cross-attention module to achieve feature-level fine-grained style incorporation. ASI consists of the proposed Siamese Cross-Attention (SiCA) and Adaptive Content-Style Blending (AdaBlending) modules, which collaborate to preserve image structure while facilitating effective style transfer.
4 Method
Given a latent representation of a real or synthetic source content image $I_{c}$ along with its corresponding content prompt $P_{c}$ and a style prompt $P_{s}$, our objective is to generate a target stylized image $I_{s}$ that maintains structural and semantic consistency with $I_{c}$ while exhibiting the stylistic characteristics described by $P_{s}$.
As discussed in Sec. 1, existing methods [19, 43, 32] commonly directly concatenate the content prompt and style prompt (e.g., a boy in anime artwork) and transform to key-value feature pair to achieve prompt-level style injection. We question this coarse-grained form of prompt-level style injection. Our motivation stems from the fact that if we do not reuse the spatial features of the reference diffusion process, and simply add a style token to the content prompt, the text-image misalignment problem in the T2I model will often result in unpredictable structural drift in the stylized image $I_{s}$. In this sense, if we explicitly parse style information that does not disrupt the structure of content features and incorporate it into content features, we can achieve more controllable results.
For this purpose, we propose the Adaptive Style Incorporation (ASI) to convert the existing cross-attention module to achieve structure-consistent feature-level fine-grained style transfer. The overall architecture of the proposed ASI is depicted in Fig. 2, which is composed of a Siamese Cross-Attention (SiCA) strategy and Adaptive Content-Style Blending (AdaBlending) module. We first introduce SiCA which employs a dual-track architecture to obtain independent content and style feature-level information (Sec. 4.1). Then, we propose the AdaBlending module to implement mask-guided fine-grained style incorporation from a structure-consistent perspective (Sec. 4.2). The proposed method can be applied to style transfer for both real and generated images without requiring training or fine-tuning, achieving a highly structurally consistent stylized image with the content image.
4.1 Siamese Cross-Attention
[…]
$F_{s}(Q,K_{s},V_{s})=\text{Softmax}(\frac{QK_{s}^{T}}{\sqrt{d}})V_{s},$ (10)
$F_{c}(Q,K_{c},V_{c})=\text{Softmax}(\frac{QK_{c}^{T}}{\sqrt{d}})V_{c},$ (11)
[…]
where $Q$ is the shared query features projected from the spatial features, $K_{s}-V_{s}$ and $K_{c}-V_{c}$ are the key-value feature pairs projected from the textual embedding of style prompt and content prompt with corresponding projection matrices, and $d$ is the channel dimension of query features.
This design aims to empower the query features, which signify the image structure, to independently attend to both the style and content prompt embeddings. Compared to prompt-level coarse-grained one-shot style injection, we can initially extract the style information from $F_{s}$ that has minimal structural impact on $F_{c}$ and adaptively incorporate it into $F_{c}$, achieving feature-level fine-grained style incorporation. We will detail it in the next section.
4.2 Adaptive Content-Style Blending
We propose a tuning-free Adaptive Content-Style Blending (AdaBlending) module to implement mask-guided fine-grained style incorporation from a structure-consistent perspective. AdaBlending first utilizes two parallel mask extractors to parse out the content-style blending region from the attention-head (Sec. 4.2.1) and spatial levels (Sec. 4.2.2), then employs the mask-guided content-style blending module (Sec. 4.2.3) to effectively stylize the target image while maintaining its structural consistency with the source content image. The overall architecture of the proposed AdaBlending is shown in Fig. 3.
Figure 3: Overall architecture of the proposed Adaptive Content-Style Blending (AdaBlending) module. AdaBlending achieves mask-guided fine-grained style incorporation through the proposed attention-head level and spatial level mask extractors, followed by the content-style blending operation in the structure-consistent manner.
4.2.1 Mask Extractor on Attention-Head Level
Many pioneering works  [24, 20, 44, 33, 57, 31, 30, 25, 53] have demonstrated that the spatial distribution of features effectively represents the style of an image. Therefore, if an attention head does not exhibit significant distribution differences in $F_{s}$ and $F_{c}$, we consider incorporating style information into this attention head is ineffective and disrupts the image structure of the target image.
[…]
$Cov_{s}={\frac{1}{m-1}}({(F_{s}^{i})^{\top}F_{s}^{i}-\frac{1}{m}{({\textbf{1}}^{\top}F_{s}^{i}})^{\top}{({\textbf{1}}^{\top}F_{s}^{i}})}),$ (12)
$Cov_{c}={\frac{1}{m-1}}({(F_{c}^{i})^{\top}F_{c}^{i}-\frac{1}{m}{({\textbf{1}}^{\top}F_{c}^{i}})^{\top}{({\textbf{1}}^{\top}F_{c}^{i}})}),$ (13)
$\ell^{i}=\frac{||Cov_{s}-Cov_{c}||^{2}_{F}}{4d^{2}},$ (14)
[…]
where $Cov_{s}$ and $Cov_{c}$ denote the spatial level covariance of style and content features, respectively. ${\|\cdot\|}^{2}_{F}$ represents the squared matrix Frobenius norm, 1 is a vector with all elements equal to 1, $m$, $d$, $i$ are flattened spatial dimension, channel dimension, and attention-head index, respectively. Then, we can obtain the attention-head mask $M_{head}\in\mathbb{R}^{h\times m\times d}$ as follows,
$M_{head}^{i}=\begin{cases}\textbf{1},\quad&\text{if}~{}~{}~{}~{}\mathop{\arg\max}_{i}\ \ (\ell^{i},n)\\ 0,\quad&\text{otherwise}\\ \end{cases}\,$ (15)
where $\mathop{\arg\max}_{i}~{}(\ell^{i},n)$ denotes the return of the index of top-$n$ attention heads with the greatest covariance distances, and $n$ is set to 6 for Stable Diffusion v1 [36]. $h$ denotes the number of heads of multi-head cross-attention.
4.2.2 Mask Extractor on Spatial Level
We further consider how to maintain structure consistency from the spatial level. Many works [56, 38, 28, 27] have demonstrated that regions with high activation correspond to crucial structural information within images. Therefore, we propose a mask extractor on the spatial level to maintain core structure information of $F_{c}$. The extractor selects criteria points of peak regions using the global max pooling of $F_{c}$, and the criteria points are denoted as $P\in\mathbb{R}^{h\times 1\times d}$. We adopt a non-learnable setting for the extractor and define regions in $F_{c}$ exceeding $\tau=\alpha\cdot P$ as crucial structural regions. $\alpha$ is a constant value and set to $0.7$. To preserve the features of these highly activated regions, we formulate the spatial mask $M_{spatial}\in\mathbb{R}^{h\times m\times d}$ at each attention-head as follows,
$M_{spatial}^{i}=\begin{cases}0,\quad&\text{if}~{}~{}~{}~{}F_{c}^{i}>\tau\\ \textbf{1},\quad&\text{otherwise}\\ \end{cases}\,$ (16)
4.2.3 Mask-Guided Content-Style Blending
With the above masks, we restrict the content features $F_{c}$ to incorporate style features $F_{s}$ from a structural-consistent perspective. We first perform logical OR operations on two sets of masks $M_{head}$ and $M_{spatial}$ to obtain the final mask $M_{all}$. Then, we perform the mask-guided blending between the structured style features powered by adaptive instance normalization operation [24], and content features to obtain the fused fine-grained stylized result as AdaBlending’s output $F_{out}$,
$F^{i}_{out}=(\sigma(F_{s}^{i})(\frac{F_{c}^{i}-\mu(F_{c}^{i})}{\sigma(F_{c}^{i})})+\mu(F_{s}^{i}))\odot M_{all}^{i}+F_{c}^{i}\odot(\textbf{1}-M_{all}^{i}),$ (17)
where $\odot$ denotes the inner product operation.
5 Experiments
In this section, we perform both qualitative and quantitative evaluations across three applications: style transfer for real-world images, visual enhancement for real-world images, and style transfer for generated images. We compare our method against state-of-the-art (SOTA) tuning-free text-guided style transfer method based on diffusion models, including PnP [43], P2P [19], and SDEdit [32], across several common style transfer tasks. Finally, we ablate each component of our method and show its effectiveness. We evaluate our method using Stable Diffusion v1 architecture [36]. Please refer to the Appendix for the implementation details of our method, quantitative evaluation results, and more quantitative results.
5.1 Qualitative Comparison
5.1.1 Style Transfer for Real-World Images
Figure 4: Comparisons with different methods on style transfer for real-world images. Sample results are shown for each of the two benchmarks: ImageNet-R-TI2I and Wild-TI2I. From left to right: the source image and text prompt, our results, P2P [19], PnP [43], SDEdit [32] with two different denoising strengths. (Zoom in for best view.)
To verify the effectiveness and generalization of ASI, we show various style transfer results applied by our approach on Wild-TI2I [43] and ImageNet-R-TI2I [18, 43] datasets, which are collected from real-world. For each result, we show images before and after stylized. As shown in Fig. 4, we compare our method against SOTA text-guided style image editing models based on diffusion models, including PnP [43], P2P [19], and SDEdit [32]. Since our method aims to incorporate style information at the fine-grained feature level, compared to prompt-level style injection, it allows us to stylize in the smaller local region of images, thereby reducing the distortion of image structure. Although SDEdit is simple but effective, it requires manual adjustment of the trade-off between the two. PNP [43] focuses more on the structural consistency of stylized images but lacks effective stylization in some cases (e.g. Row. 3 and 5). While P2P [19] is the opposite, effective stylization comes at the expense of structural and instance deviation (e.g. stone in the Row. 3). As shown in Col. 2 of Fig. 4, our method effectively stylizes images while maintaining structural consistency with the source image, achieving a good balance between stylization and structural preservation compared to other SOTA methods.
Figure 5: Comparisons with different methods on visual enhancement for real-world images. From left to right: the source image, visual enhancement prompts, our results, P2P [19], PnP [43], SDEdit [32]. (Zoom in for best view.)
Figure 6: Comparisons with different methods on style transfer for generated images. From left to right: the source image and text prompt, our results, P2P [19], PnP [43], SDEdit [32] with two different denoising strengths. (Zoom in for best view.)
5.1.2 Visual Enhancement for Real-World Images
Benefiting from the text-guided style transfer manner instead of the image-guided one, the diffusion-based method can be broadened to application areas unreachable by traditional image-to-image style transfer methods [58, 24, 14], particularly those related to visual enhancements such as vividness, realism, and depth of field. This is because these "styles" cannot be represented through an image.
As shown in Fig. 5, while P2P [19], PNP [43], and SDEdit [32] are based on the diffusion model, they only preserve the overall structure of the source image, sacrificing the detailed structure information for achieving effective visual enhance. We only present SDEdit with $0.6$ denoising strength, because excessive denoising strength usually causes drastic drifts in the image structure. This limits the application of the above methods to this specific task, as it necessitates strict consistency between the source and target images. Thanks to our proposed fine-grained style incorporation idea, our method efficiently enhances the visual quality of images while maintaining structural consistency. In terms of appearance enhancement, our method better maintains the horse’s body (Row. 1), the cat’s eyes (Row. 2), and the goldfish’s body details (Row. 3). In addition, our method achieves blurring under scene-invariant and semantic-invariant conditions, such as Spiderman’s left arm (Row. 4), and background details in Row. 5 and 6. These impressive results reveal the immense potential of diffusion models in addressing low-level vision tasks.
5.1.3 Style Transfer for Generated Images
As shown in Fig. 6, our method effectively transfers the source image to the target style while maintaining a high degree of structural consistency with the source image. Compared to our method, other methods are more prone to exhibiting structure deviations. As shown in Row. 1-3, the results of other methods can preserve the overall structure of the source image, yet they introduce notable modifications to the key visual features of the object. Furthermore, in Row. 4-6, the stylized images fail to maintain overall structural consistency with the source image. Our results achieve effective style transfer while minimizing modifications to the source image.
Figure 7: Ablation study on the guidance mask of the AdaBlending module.
5.2 Ablation Study
We conduct ablation experiments on the guidance mask of the AdaBlending module. As shown in Fig. 7, when both spatial level and attention-head level masks are simultaneously utilized (Col. 2), AdaBlending achieves consistent structural and semantic style transfer results. However, when only one guidance mask is used (Col. 3 and 4), noticeable structural drifts are observed in the girl’s head and eyebrow details. Furthermore, without using any guidance mask (Col. 5), additional semantic drifts are observed in the T-shirt color. These results highlight the importance of guidance masks in ensuring accurate and consistent style transfer outcomes.
6 Conclusion
In this work, we aim to study text-driven style transfer. To mitigate the well-known issue of structure inconsistency induced by text-image misalignment in diffusion-based T2I models, we propose a novel and effective tuning-free algorithm, termed Adaptive Style Incorporation (ASI), to incorporate the style information on the fine-grained feature level. It consists of the Siamese Cross-Attention (SiCA) to decouple the single-track cross-attention to a dual-track structure to obtain separate content and style features and the Adaptive Content-Style Blending (AdaBlending) module to couple the content and style information from a structure-consistent manner. We conduct a thorough evaluation of our method across diverse source image domains, both real and generated ones, on a range of common style transfer tasks. Experimental results have demonstrated the effectiveness of our proposed ASI and exhibit significant advantages over SOTA methods. Limitations and future work are provided in the Appendix.